What is the molarity of 0.349 g NaCl in 0.7 L of water? Use 58.4 for the molecular weight of NaCl.
Answer #1

Molarity of NaCl = Moles of solutes/volume of solution in liters

In our case, solute is NaCl.

Molar mass of NaCl= 23 + 35.4 = 58.5 g/mol.

Number of moles = given mass/molar mass

We are given the mass of NaCl, i.e., 0.349 g

Moles of NaCl = 0.349/58.4

=0.00597

Volume of solution = 0.7 L

Molarity of solution = 0.00597/0.7

= 0.008 M